OpenStudy (anonymous):

Can somebody please show me how this equation is done. I will give a medal and become a fan. x-2y=-8

OpenStudy (anonymous):

well this equation have infinite number of solution points.

OpenStudy (anonymous):

if you choose y = 0, then x = -8, so (-8,0) is one of the solution of the equation

OpenStudy (anonymous):

i have to turn it into slope intercept form and graph @sourav_aich

OpenStudy (anonymous):

the slope of the equation can be found by the following steps: the given equation is x - 2y = -8 => -2y = -x - 8 => y = 0.5x + 4 comparing the above equation with the slope intercept form of st. line y = mx + c where 'm' is the slope and 'c' is the y intercept we can say that the slope of the given line is 0.5 and y intercept is 4